OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through (-3, 5) and is parallel to the graph of y = 2x - 4.

OpenStudy (anonymous):

slope of line will be 2 as the coefficient of x is 2. given point is(-3,5) so required equation, (y-y1)=m(x-x1) (y-5)=2(x+3)

OpenStudy (loser66):

HELLO, DUDE..IM..A..DUCK!! IM.. A..PENGUIN hehehe

OpenStudy (anonymous):

Your not a duck

OpenStudy (anonymous):

WELL SCREW YOU TOO

OpenStudy (anonymous):

not a duck pshhh im obviosly a duck

OpenStudy (anonymous):

All i said is your not a duck

OpenStudy (anonymous):

that hurts my feelings ;-;

OpenStudy (anonymous):

So Quack Quack to you

OpenStudy (anonymous):

incog tell them they are rong

OpenStudy (anonymous):

wrooooooongggg

OpenStudy (lncognlto):

Substitute m = 2 and (-3,5) in y - y1 = m(x - x1): therefore y - 5 = 2(x - (-3)) tf y - 5 = 2x + 6 tf y = 2x + 11... XD

OpenStudy (anonymous):

nope

OpenStudy (lncognlto):

They are most definitely wrong! Duck is a duck!

OpenStudy (lncognlto):

An evil duck, at that. xD

OpenStudy (anonymous):